OpenStudy (brittany):

solve using the quadtratic formula 2x^2=4-7x

OpenStudy (anonymous):

Put into standard form, ax^2 + bx + c = 0 2x^2 + 7x - 4 = 0 a = 2 b = 7 c = -4 Just plug it into the quadratic formula. x = -4, 1/2

OpenStudy (anonymous):

\[[-7\pm \sqrt{7^2-4*2*(-4)}]/2*2\]

OpenStudy (anonymous):

\[(-7\pm \sqrt{81})/4\]

OpenStudy (anonymous):

2x^2+7x-4=0 (2x-1)(x+4)=0 x=1/2,-4